# UNeMo: Collaborative Visual-Language Reasoning and Navigation via a Multimodal World Model

### v1

@@ Introduction
Vision-and-Language Navigation (VLN) requires an embodied agent to achieve autonomous navigation from the current to the target position in an unknown environment based on visual observations and language instructions (anderson2018vision_r2r; krantz2020beyond_vlnce; qi2020reverie). This capability constitutes the core foundation of an embodied intelligent system. This task not only needs to address the problem of efficient fusion of cross-modal information (gu2022vision_svy; wu2024vision_svy) but also poses new challenges to the construction of a perception-reasoning collaborative mechanism in dynamic environments (zhang2024vision_svy).
Traditional VLN methods use tailored end-to-end deep learning frameworks, such as sequence-to-sequence models (nips/FriedHCRAMBSKD18), attention-based cross-modal fusion (cvpr/Hong0QOG21), and graph-based representation learning (eccv/KrantzWMBL20). These jointly optimize vision-language alignment and action policy learning to map multimodal inputs to navigation decisions (an2024etpnav; yue2024safe). While performing well on large task-specific datasets, they show poor policy generalization in out-of-distribution environments or with novel instructions.
@@
Large language models (LLMs) have recently shown strong abilities in knowledge generalization and complex reasoning, prompting research into their use for enhancing VLN with promising results. Researchers have attempted to use LLMs to assist VLN in tasks such as navigation instruction parsing (yang2024llm), enhancing scene descriptions (wu2024voronav), and extracting task knowledge (shah2023lm; schumann2024velma; wang2023voyager). NavGPT further leverages the decision-making capabilities of LLMs to directly predict navigation actions (aaai/ZhouHW24). However, due to the lack of training on navigation data, the performance of such methods is inferior to traditional VLN-specialized approaches.
Accordingly, NavGPT2 employs navigation data to conduct knowledge distillation on LLMs, thereby constructing a navigation-specialized LLM (zhou2024navgpt2), as illustrated in Fig. 1. Equipped with a navigational reasoning LLM, this approach achieves performance on par with VLN-specialized methods. However, NavGPT2 is trained exclusively via language-guided reasoning modules, rendering it incapable of visual state reasoning. Furthermore, its navigational reasoning module is frozen during the training phase of navigation policies, creating a disconnect between the reasoning process and navigation optimization. This hinders the dynamic refinement of navigational reasoning capabilities in response to policy training progress, thereby imposing inherent limitations on navigation performance.
@@
Traditional research in Vision-Language Navigation (VLN) has centered on state representation learning, including recurrent neural network (RNN)-based historical state modeling  (Hong_2021), sequential encoding for temporal information processing  (chen2021history), topological map construction for spatial connectivity  (an2024etpnav; chen2022think_topo), and grid map development for precise localization  (liu2023bird_map), Additionally, frameworks such as SUSA  (zhang2024agent_SUSA) have introduced hybrid semantic-spatial representations to enhance vision-language alignment. Recent advances have shifted toward leveraging Large Language Models (LLMs) to enhance navigation capabilities.
Navigation with LLM. LLMs, with their strong semantic understanding and reasoning abilities, have been increasingly applied to navigation agents (huang2022inner; yao2023react; shah2023lm; schumann2024velma; wang2023voyager). Early approaches decoupled LLMs from core navigation systems—using GPT-2 for subtask hints (qiao2023march) or GPT-4 prompts to translate instructions into actions (song2025towards)—limiting holistic integration. Recent work embeds LLMs directly, such as zero-shot NavGPT (aaai/ZhouHW24), MapGPT (chen2024mapgpt), MC-GPT (zhan2024mc), and fine-tuned LLaMA-7B variants like LangNav (pan2023langnav) and NavCoT (lin2025navcot). Despite improvements, these agents still lag behind specialized VLN models in accuracy and transparency. NavGPT-2, which freezes the LLM as a vision-language encoder with imitation learning (zhou2024navgpt2), lacks dynamic scene adaptation. In contrast, our work integrates a future-state prediction module to dynamically forecast environmental changes for enhanced decision-making.
World Model for VLN. World models (Ha_Schmidhuber_2018) compress high-dimensional sensory data into dynamic state representations, supporting embodied decision-making. Prior studies explored diverse architectures for environmental feature encoding  (Chang_2020; Purushwalkam_2020), while sampling-based planners used these representations for self-supervised policy learning (Sekar_2020; Chen20a). Pathdreamer (Koh_2021_ICCV) and Dreamwalker (wang2023dreamwalker) predict future observations (pixel-level or abstract) to extend planning horizons but focus solely on forward exploration. LookAhead  (wang2024lookahead) employs a hierarchical neural radiance field to generate multi-scale future embeddings. However, these methods separate state prediction from policy learning. Our work introduces a joint framework for simultaneous future state prediction and proposes a hierarchical prediction-feedback mechanism that co-optimizes state inference and action selection, fully integrating predicted future states into navigation decisions for more precise navigation.
@@
The action prediction in VLN can be formulated as a node-selection problem. Early methods confined this space to the local neighborhood of the agent’s current node; more recent approaches have extended it to the global scale by constructing an explicit topological map (chen2022think_topo; zhou2024navgpt2). The agent builds a topological map $\mathcal{G}=(\mathcal{V},\mathcal{E})$ based on the nodes it has visited and the nodes within its visual range, where $\mathcal{V}=\{v_{1},v_{2},...,v_{N}\}$ denotes the set of N navigable nodes, and $\mathcal{E}$ represents the set of connecting edges between nodes. At each time step $t$, the agent acquires observations of adjacent navigable nodes at the current position and updates the topological map $\mathcal{G}_{t}$. For visited nodes, the node embedding is the observation at that node. For unvisited but partially observable nodes, the node embedding aggregates the average of all partial-view embeddings observed from other visited nodes. The final action prediction $a_{t}$ corresponds to selecting a node $v\in\mathcal{V}$.
@@
We adopt NavGPT2 (zhou2024navgpt2), one of the most advanced LLM-based navigation approaches to date, as our base model. NavGPT2 incorporates three core components: node embedding, decision making, and loss function. For node embedding, it employs a Topology Node Embedding (TNE) encoder. Specifically, let $E^{o}$ and $E^{x}$ denote the visual (observation) and textual (instruction) features, respectively. These are jointly encoded by a pretrained LLM to produce the corresponding features $F^{o}$ and $F^{x}$. The TNE encoder then projects these embeddings into a shared topological space to form the node embeddings used for navigation:
@@
For the loss function, NavGPT2 integrates behavioral cloning and DAgger (ross2011reduction) objectives into a single training criterion:
@@
While topological maps encode historical states—encompassing visual observations and node connectivity—they lack information about future states. Given that future state prediction enhances navigation performance (li2023improving; wang2023dreamwalker; wang2024lookahead), we introduce a Multimodal World Model (MWM) to predict future visual states, leveraging the topological map’s representation of past states. We employ Conditional Variational Autoencoders (CVAE) (NIPS2015_cvae) to construct the MWM, which comprises an encoder for aligning and fusing multimodal inputs, and a decoder for reasoning about the next-step visual state.
@@
Our work primarily focus on the path instruction following task on the Room-to-Room (R2R) benchmark dataset  (anderson2018vision_r2r). R2R requires an agent to follow step-by-step natural language instructions, averaging 32 words in length. To rigorously evaluate the generalization capabilities of our approach, we additionally conduct evaluations on the REVERIE benchmark  (qi2020reverie). REVERIE presents more challenging high-level instructions (averaging 21 words), such as ”Find the cell phone on the nightstand in the bedroom”, and requires the agent to navigate to the target location and localize a distant target object within a pre-defined ground-truth bounding box. Notably, the expert demonstration path lengths in REVERIE range from 4 to 7 steps, compared to an average of 6 steps in R2R.
@@
We adopt the widely recognized evaluation framework in vision-and-language navigation. For core navigation performance assessment, key metrics include: Success Rate (SR), measuring the proportion of paths where the agent’s final position is within a preset threshold (typically 3 meters) from the goal; Success weighted by Path Length (SPL), which jointly considers task success and path efficiency; Trajectory Length (TL) representing the average path length in meters; Navigation Error (NE) indicating the average distance between the agent’s final position and the target; and Oracle Success Rate (OSR) showing success rate under ideal stopping policy. For object referral evaluation, following REVERIE (qi2020reverie), we employ Remote Grounding Success (RGS) to assess the proportion of successfully reaching the correct area and accurately locating the target object, along with its path-length-weighted variant (RGSPL). Among all metrics, except TL and NE where lower values indicate better performance, higher values represent superior results for other indicators.
@@
To demonstrate that UNeMo serves as a generalizable framework applicable to diverse baseline architectures, we validate its effectiveness on two distinct topology-based navigation policies: NavGPT2 (zhou2024navgpt2) and DUET (chen2022think_topo). Both policies share a common limitation: while capable of processing vision-language inputs and constructing topological representations via imitation learning, they lack prospective state reasoning capabilities. To address this gap, we integrate explicit state reasoning tasks into both systems using the UNeMo framework. Our experiments retain all baseline pretraining results and introduce UNeMo exclusively during the fine-tuning phase. An important implementation distinction lies in their modality processing: DUET employs conventional vision-language pretraining for feature extraction, whereas NavGPT2 acquires representations via LLM intermediate layer distillation. This difference provides a controlled setting for evaluating UNeMo’s adaptability to diverse multimodal inputs. All experiments were conducted on Ubuntu 22.04 using NVIDIA RTX 4090 GPUs (single-GPU configuration), with complete environment specifications available in our open-source code repository
@@
Table 2: Performance Comparison on REVERIE Dataset.*indicates reproduced results
@@
For the R2R dataset experiments, we compared our method with state-of-the-art (SOTA) approaches in LLM-based vision-and-language navigation, including NavCoT (lin2025navcot), LangNav(pan2023langnav), NaviLLM(zheng2024NaviLLM), and NavGPT2(zhou2024navgpt2). These methods formulate navigation as a language modeling problem, employing either prompt engineering or simple imitation learning paradigms for end-to-end navigation decision-making.
@@
Table 3: Comparison of Different State Reasoning Methods. “TopoState” denotes the topology-map-based approach; “Cond2Vis” denotes vision-based approach; “Vis-WM” denotes the vision-based world model; “UNeMo” denotes our multi-modal-based world model.
@@ Acknowledgements
Prior to finalizing the multimodal world model as our ultimate visual state reasoning approach, we systematically investigated various alternative solutions (as detailed in the ablation studies section). These exploratory endeavors laid crucial groundwork for the development of the UNeMo method and provided significant insights into the pivotal role of state reasoning in agent navigation. This section elaborates on the design and implementation of these methods, followed by a comparative analysis of their performance on the R2R dataset (anderson2018vision_r2r).
Topology map-based state reasoning approach.
Inspired by the topological mapping mechanism in NavGPT2 (zhou2024navgpt2), we identified two critical phases in its navigation pipeline: (1) updating node encodings in the topological map based on current visual observations and natural language instructions upon arriving at each node, and (2) scoring globally navigable nodes according to the node encodings in the topological map (see the Preliminary section for details). Building upon this framework, we designed a supervised learning task to enhance the agent’s state reasoning capability. Specifically, we first aggregate all node features $\mathcal{V}_{t}=\{\mathcal{V}_{t,i}\}_{i=1}^{K}$ in the current topological map $\mathcal{G}_{t}$ (where $K$ denotes the number of nodes), via average pooling to obtain the current state representation $S^{m}_{t}$:
@@
Vision-Based Decoder Schemes.
@@
Dense Visual Prediction: Utilizes the current topological node encoding $\mathcal{V}_{t}$ to predict the dense representation of visual observations at candidate nodes. The target representation is obtained via Vision Transformer (ViT) from next-state visual observations $O_{t+1}$:
@@
Sparse Visual Prediction: Employs a Conditional Variational Autoencoder (CVAE) to predict compressed representations of post-action visual observations from partial views $O_{t}$. First, an autoencoder compresses the full visual representation of ground-truth nodes:
@@
Baseline implementation differences. The multimodal state reasoning module in the UNeMo framework requires both visual-linguistic modality data as input and incorporates topological map node encodings as conditional information. Experimental validation was conducted on two topology-based baseline methods: the NavGPT2-based UNeMo implementation (detailed in the Methods section) and the DUET-based modified version. The primary distinction lies in their multimodal feature extraction approaches - the former employs frozen latent representations from a large language model (LLM) for visual-linguistic encoding, whereas the latter utilizes representations generated by trainable Transformer blocks.
Training strategy details. To prevent overfitting in the multimodal state prediction module and enhance multi-task training stability during navigation tasks, we adopt a phased training strategy. Specifically, the training process is divided into multiple phases where state reasoning tasks are only activated during the initial 10% of training batches in each phase. This design addresses two critical considerations: first, while multimodal state reasoning primarily aims to improve navigation performance by enhancing the agent’s understanding of sequential navigation processes, direct joint training would cause premature convergence of the CVAE module due to insufficient state reasoning data, thereby compromising the learning efficacy of the navigation policy network; second, since conditional information dynamically evolves with the training of the navigation policy network, we periodically retrain the Multimodal World Model (MWM) to maintain alignment with the shifting feature space.

### v2

@@ Introduction
Vision-and-Language Navigation (VLN) requires an embodied agent to achieve autonomous navigation from the current to the target position in an unknown environment based on visual observations and language instructions (Anderson et al. 2018; Krantz et al. 2020b; Qi et al. 2020). This capability constitutes the core foundation of an embodied intelligent system. This task not only needs to address the problem of efficient fusion of cross-modal information (Gu et al. 2022; Wu et al. 2024b) but also poses new challenges to the construction of a perception-reasoning collaborative mechanism in dynamic environments (Zhang et al. 2024b).
Traditional VLN methods use tailored end-to-end deep learning frameworks, such as sequence-to-sequence models (Fried et al. 2018), attention-based cross-modal fusion (Hong et al. 2021a), and graph-based representation learning (Krantz et al. 2020a). These jointly optimize vision-language alignment and action policy learning to map multimodal inputs to navigation decisions (An et al. 2024; Yue et al. 2024). While performing well on large task-specific datasets, they show poor policy generalization in out-of-distribution environments or with novel instructions.
@@
Large language models (LLMs) have recently shown strong abilities in knowledge generalization and complex reasoning, prompting research into their use for enhancing VLN with promising results. Researchers have attempted to use LLMs to assist VLN in tasks such as navigation instruction parsing (Yang et al. 2024), enhancing scene descriptions (Wu et al. 2024a), and extracting task knowledge (Shah et al. 2023; Schumann et al. 2024; Wang et al. 2023a). NavGPT further leverages the decision-making capabilities of LLMs to directly predict navigation actions (Zhou et al. 2024b). However, due to the lack of training on navigation data, the performance of such methods is inferior to traditional VLN-specialized approaches.
Accordingly, NavGPT2 employs navigation data to conduct knowledge distillation on LLMs, thereby constructing a navigation-specialized LLM (Zhou et al. 2024a), as illustrated in Fig. 1. Equipped with a navigational reasoning LLM, this approach achieves performance on par with VLN-specialized methods. However, NavGPT2 is trained exclusively via language-guided reasoning modules, rendering it incapable of visual state reasoning. Furthermore, its navigational reasoning module is frozen during the training phase of navigation policies, creating a disconnect between the reasoning process and navigation optimization. This hinders the dynamic refinement of navigational reasoning capabilities in response to policy training progress, thereby imposing inherent limitations on navigation performance.
@@
Traditional research in Vision-Language Navigation (VLN) has centered on state representation learning, including recurrent neural network (RNN)-based historical state modeling  (Hong et al. 2021b), sequential encoding for temporal information processing  (Chen et al. 2021), topological map construction for spatial connectivity  (An et al. 2024; Chen et al. 2022), and grid map development for precise localization  (Liu et al. 2023), Additionally, frameworks such as SUSA  (Zhang et al. 2024a) have introduced hybrid semantic-spatial representations to enhance vision-language alignment. Recent advances have shifted toward leveraging Large Language Models (LLMs) to enhance navigation capabilities.
Navigation with LLM. LLMs, with their strong semantic understanding and reasoning abilities, have been increasingly applied to navigation agents (Huang et al. 2022; Yao et al. 2023; Shah et al. 2023; Schumann et al. 2024; Wang et al. 2023a). Early approaches decoupled LLMs from core navigation systems—using GPT-2 for subtask hints (Qiao et al. 2023) or GPT-4 prompts to translate instructions into actions (Song et al. 2025)—limiting holistic integration. Recent work embeds LLMs directly, such as zero-shot NavGPT (Zhou et al. 2024b), MapGPT (Chen et al. 2024), MC-GPT (Zhan et al. 2024), and fine-tuned LLaMA-7B variants like LangNav (Pan et al. 2023) and NavCoT (Lin et al. 2025). Despite improvements, these agents still lag behind specialized VLN models in accuracy and transparency. NavGPT-2, which freezes the LLM as a vision-language encoder with imitation learning (Zhou et al. 2024a), lacks dynamic scene adaptation. In contrast, our work integrates a future-state prediction module to dynamically forecast environmental changes for enhanced decision-making.
World Model for VLN. World models (Ha and Schmidhuber 2018) compress high-dimensional sensory data into dynamic state representations, supporting embodied decision-making. Prior studies explored diverse architectures for environmental feature encoding  (Chang et al. 2020; Purushwalkam et al. 2020), while sampling-based planners used these representations for self-supervised policy learning (Sekar et al. 2020; Yen-Chen et al. 2020). Pathdreamer (Koh et al. 2021) and Dreamwalker (Wang et al. 2023b) predict future observations (pixel-level or abstract) to extend planning horizons but focus solely on forward exploration. LookAhead  (Wang et al. 2024) employs a hierarchical neural radiance field to generate multi-scale future embeddings. However, these methods separate state prediction from policy learning. Our work introduces a joint framework for simultaneous future state prediction and proposes a hierarchical prediction-feedback mechanism that co-optimizes state inference and action selection, fully integrating predicted future states into navigation decisions for more precise navigation.
@@
The action prediction in VLN can be formulated as a node-selection problem. Early methods confined this space to the local neighborhood of the agent’s current node; more recent approaches have extended it to the global scale by constructing an explicit topological map (Chen et al. 2022; Zhou et al. 2024a). The agent builds a topological map $\mathcal{G}=(\mathcal{V},\mathcal{E})$ based on the nodes it has visited and the nodes within its visual range, where $\mathcal{V}=\{v_{1},v_{2},...,v_{N}\}$ denotes the set of N navigable nodes, and $\mathcal{E}$ represents the set of connecting edges between nodes. At each time step $t$, the agent acquires observations of adjacent navigable nodes at the current position and updates the topological map $\mathcal{G}_{t}$. For visited nodes, the node embedding is the observation at that node. For unvisited but partially observable nodes, the node embedding aggregates the average of all partial-view embeddings observed from other visited nodes. The final action prediction $a_{t}$ corresponds to selecting a node $v\in\mathcal{V}$.
@@
We adopt NavGPT2 (Zhou et al. 2024a), one of the most advanced LLM-based navigation approaches to date, as our base model. NavGPT2 incorporates three core components: node embedding, decision making, and loss function. For node embedding, it employs a Topology Node Embedding (TNE) encoder. Specifically, let $E^{o}$ and $E^{x}$ denote the visual (observation) and textual (instruction) features, respectively. These are jointly encoded by a pretrained LLM to produce the corresponding features $F^{o}$ and $F^{x}$. The TNE encoder then projects these embeddings into a shared topological space to form the node embeddings used for navigation:
@@
For the loss function, NavGPT2 integrates behavioral cloning and DAgger (Ross et al. 2011) objectives into a single training criterion:
@@
While topological maps encode historical states—encompassing visual observations and node connectivity—they lack information about future states. Given that future state prediction enhances navigation performance (Li and Bansal 2023; Wang et al. 2023b, 2024), we introduce a Multimodal World Model (MWM) to predict future visual states, leveraging the topological map’s representation of past states. We employ Conditional Variational Autoencoders (CVAE) (Sohn et al. 2015) to construct the MWM, which comprises an encoder for aligning and fusing multimodal inputs, and a decoder for reasoning about the next-step visual state.
@@
Our work primarily focus on the path instruction following task on the Room-to-Room (R2R) benchmark dataset  (Anderson et al. 2018). R2R requires an agent to follow step-by-step natural language instructions, averaging 32 words in length. To rigorously evaluate the generalization capabilities of our approach, we additionally conduct evaluations on the REVERIE benchmark  (Qi et al. 2020). REVERIE presents more challenging high-level instructions (averaging 21 words), such as ”Find the cell phone on the nightstand in the bedroom”, and requires the agent to navigate to the target location and localize a distant target object within a pre-defined ground-truth bounding box. Notably, the expert demonstration path lengths in REVERIE range from 4 to 7 steps, compared to an average of 6 steps in R2R.
@@
We adopt the widely recognized evaluation framework in vision-and-language navigation. For core navigation performance assessment, key metrics include: Success Rate (SR), measuring the proportion of paths where the agent’s final position is within a preset threshold (typically 3 meters) from the goal; Success weighted by Path Length (SPL), which jointly considers task success and path efficiency; Trajectory Length (TL) representing the average path length in meters; Navigation Error (NE) indicating the average distance between the agent’s final position and the target; and Oracle Success Rate (OSR) showing success rate under ideal stopping policy. For object referral evaluation, following REVERIE (Qi et al. 2020), we employ Remote Grounding Success (RGS) to assess the proportion of successfully reaching the correct area and accurately locating the target object, along with its path-length-weighted variant (RGSPL). Among all metrics, except TL and NE where lower values indicate better performance, higher values represent superior results for other indicators.
@@
To demonstrate that UNeMo serves as a generalizable framework applicable to diverse baseline architectures, we validate its effectiveness on two distinct topology-based navigation policies: NavGPT2 (Zhou et al. 2024a) and DUET (Chen et al. 2022). Both policies share a common limitation: while capable of processing vision-language inputs and constructing topological representations via imitation learning, they lack prospective state reasoning capabilities. To address this gap, we integrate explicit state reasoning tasks into both systems using the UNeMo framework. Our experiments retain all baseline pretraining results and introduce UNeMo exclusively during the fine-tuning phase. An important implementation distinction lies in their modality processing: DUET employs conventional vision-language pretraining for feature extraction, whereas NavGPT2 acquires representations via LLM intermediate layer distillation. This difference provides a controlled setting for evaluating UNeMo’s adaptability to diverse multimodal inputs. All experiments were conducted on Ubuntu 22.04 using NVIDIA RTX 4090 GPUs (single-GPU configuration), with complete environment specifications to be provided in our open-source code repository later.
@@
Table 2: Performance Comparison on REVERIE Dataset. *indicates reproduced results.
@@
For the R2R dataset experiments, we compared our method with state-of-the-art (SOTA) approaches in LLM-based vision-and-language navigation, including NavCoT (Lin et al. 2025), LangNav(Pan et al. 2023), NaviLLM(Zheng et al. 2024), and NavGPT2(Zhou et al. 2024a). These methods formulate navigation as a language modeling problem, employing either prompt engineering or simple imitation learning paradigms for end-to-end navigation decision-making.
@@
Table 3: Comparison of Different State Reasoning Methods. “TopoState” denotes the topology-map-based approach, “Cond2Vis” denotes vision-based approach, “Vis-WM” denotes the vision-based world model, “UNeMo” denotes our multi-modal-based world model.
@@ Acknowledgements
Prior to finalizing the multimodal world model as our ultimate visual state reasoning approach, we systematically investigated various alternative solutions (as detailed in the ablation studies section). These exploratory endeavors laid crucial groundwork for the development of the UNeMo method and provided significant insights into the pivotal role of state reasoning in agent navigation. This section elaborates on the design and implementation of these methods, followed by a comparative analysis of their performance on the R2R dataset (Anderson et al. 2018).
Topology map-based state reasoning approach
Inspired by the topological mapping mechanism in NavGPT2 (Zhou et al. 2024a), we identified two critical phases in its navigation pipeline: (1) updating node encodings in the topological map based on current visual observations and natural language instructions upon arriving at each node, and (2) scoring globally navigable nodes according to the node encodings in the topological map (see the Preliminary section for details). Building upon this framework, we designed a supervised learning task to enhance the agent’s state reasoning capability. Specifically, we first aggregate all node features $\mathcal{V}_{t}=\{\mathcal{V}_{t,i}\}_{i=1}^{K}$ in the current topological map $\mathcal{G}_{t}$ (where $K$ denotes the number of nodes), via average pooling to obtain the current state representation $S^{m}_{t}$:
@@
Vision-Based Decoder Schemes
@@
Dense Visual Prediction
Utilizes the current topological node encoding $\mathcal{V}_{t}$ to predict the dense representation of visual observations at candidate nodes. The target representation is obtained via Vision Transformer (ViT) from next-state visual observations $O_{t+1}$:
@@
Sparse Visual Prediction
Employs a Conditional Variational Autoencoder (CVAE) to predict compressed representations of post-action visual observations from partial views $O_{t}$. First, an autoencoder compresses the full visual representation of ground-truth nodes:
@@
aseline implementation differences
The multimodal state reasoning module in the UNeMo framework requires both visual-linguistic modality data as input and incorporates topological map node encodings as conditional information. Experimental validation was conducted on two topology-based baseline methods: the NavGPT2-based UNeMo implementation (detailed in the Methods section) and the DUET-based modified version. The primary distinction lies in their multimodal feature extraction approaches - the former employs frozen latent representations from a large language model (LLM) for visual-linguistic encoding, whereas the latter utilizes representations generated by trainable Transformer blocks.
Table 5: Performance Comparison on R2R Dataset
Method Freeze LLM Val Unseen Test Unseen
$uparrow SR\uparrow SPL\uparrow$
NavGPT2 (1.5B) 70.0 59.0 71.0 60.0
NavGPT2 (5B) 72.0 61.0 71.0 60.0
UNeMo (1.5B) 72.1 61.1 72.5 61.3
Table 6: Generalization Performance on REVERIE Dataset
Method Val Unseen Test Unseen
$uparrow SR\uparrow SPL\uparrow$
DUET* 46.66 31.03 52.61 34.82
UNeMo 49.36 32.84 53.21 35.17
Training strategy details
To prevent overfitting in the multimodal state prediction module and enhance multi-task training stability during navigation tasks, we adopt a phased training strategy. Specifically, the training process is divided into multiple phases where state reasoning tasks are only activated during the initial 10% of training batches in each phase. This design addresses two critical considerations: first, while multimodal state reasoning primarily aims to improve navigation performance by enhancing the agent’s understanding of sequential navigation processes, direct joint training would cause premature convergence of the CVAE module due to insufficient state reasoning data, thereby compromising the learning efficacy of the navigation policy network; second, since conditional information dynamically evolves with the training of the navigation policy network, we periodically retrain the Multimodal World Model (MWM) to maintain alignment with the shifting feature space.